# Can femtoscopic correlation function shed light on the nature of the lightest, charm, axial mesons?

### v1

@@ Abstract
There seem to exist two lightest axial mesons with charm whose masses are very similar but the associated widths and other properties are different. These mesons are denominated as $D_{1}(2430)$ and $D_{1}(2420)$. Although two mesons with similar masses are expected to exist, with such quantum numbers, within the traditional quark model, as we discuss the description of their decay widths and other properties requires contributions from meson-meson coupled channel scattering. We present the amplitudes obtained by solving the Bethe-Salpeter equations which unavoidably lead to the generation of two axial resonances when mesons are considered as the degrees of freedom in the model. One of them is narrow and has properties in good agreement with those of $D_{1}(2420)$. The other pole is wider, but not wide enough to be related to $D_{1}(2430)$. Its position (mass) also does not match well with that of $D_{1}(2430)$. The situation improves when a bare quark-model state is included. Further, we calculate the scattering lengths for different channels and compare them with the available data from lattice QCD. Using arguments of heavy quark symmetry, we discuss possible differences between the scattering lengths coming from the lattice QCD calculations for the $D\pi$ and $D^{*}\pi$ systems and the preliminary experimental data from the ALICE Collaboration on the $D\pi$ system. We present two models, which can produce compatible properties for the two lightest $D_{1}$ states but result in different scattering lengths: one in agreement with the findings of lattice QCD and the other in agreement with the estimation obtained using the $D\pi$ results from the ALICE Collaboration. We present the correlation functions for both cases and discuss how femtoscopic physics can shed light on this issue.
@@ I Introduction
Femtoscopic correlation functions have emerged as a powerful alternative tool to understand the nature of hadrons in recent times (to see the articles published in the current year, for example, see Refs. Acharya et al. (2023a, b, c) or see the review of Ref. Fabbietti et al. (2021)), though the idea is not new Lisa et al. (2005). In the present manuscript, we analyze if the same idea can be implemented to determine the properties of the two lightest $D_{1}$ mesons, $D_{1}(2420)$ and $D_{1}(2430)$, for which the average values of mass and width, as given by the Particle Data Group (PDG) Workman et al. (2022), are:
@@
The evidence for the existence of two $D_{1}$ states, with the aforementioned properties, comes from a fit made to the experimental data on the $D^{*}\pi$ invariant mass distribution Aaij et al. (2015); Abe et al. (2004), where the $S$- and $D$-wave amplitudes are associated with $D_{1}(2430)$ and $D_{1}(2420)$ respectively. From the theory side, the existence of two $c\bar{d}$, $c\bar{u}$ states were predicted, with the spin-parity $J^{\pi}=1^{+}$, in a quark model based on one-gluon-exchange-plus-linear-confinement potential Godfrey and Isgur (1985). Such states were found to arise from the $1^{1}P_{1}$ and $1^{3}P_{1}$ interactions, with masses 2440 MeV and 2490 MeV, respectively, Godfrey and Isgur (1985). An admixture of the states is associated with physical mesons, with a mixing angle of $-41^{0}$ Godfrey and Isgur (1985). Though the masses are not very different from those listed in Eq. (1), the two states would be expected to have similar widths from the point of view of the quark model (as indeed found in Ref. Ferretti and Santopinto (2018), where a width of $\sim 30$ MeV is obtained for both states), thus disagreeing with the experimental findings.
@@
On the other hand, following a different formalism, it was pointed out in Ref. Du et al. (2018) that all the low-lying positive-parity heavy open-flavor mesons can be understood as hadronic molecules. Such claims are motivated by facts like the mass of the lightest open-charm mesons with strangeness are lower than their non-strange counterparts. In Ref. Du et al. (2018), scattering equations were solved using kernels coming from the unitarized chiral perturbation theory for heavy mesons and by fixing the free parameters to fit the scattering lengths determined from Lattice QCD calculations. Within such a framework, two poles were found for $D_{1}$-mesons, with mass, half-width: $2247^{+5}_{-6},~{}107^{+11}_{-10}$ and $2555^{+47}_{-30},~{}203^{+8}_{-9}$ and the lower one was associated with $D_{1}(2430)$. It is argued that the mass value, obtained from a Breit-Wigner fit, given by PDG needs to be revised. A narrow pole, to be associated with $D_{1}(2420)$, however, is not found. Curiously, a very different quark-model calculation Abreu et al. (2019) obtained mass values for the $D_{1}$ states which are very similar to those of Ref. Du et al. (2018). Information on $D_{1}$ mesons is also available from ${}^{2S+1}l_{J}={}^{3}S_{1},~{}{}^{3}D_{1}$ amplitudes determined for the $D^{*}\pi$ system with lattice QCD calculations Lang and Wilson (2022). A broad and a narrow state is found, respectively, in the former and the latter amplitudes. The broad ${}^{3}S_{1}$ state, which is related to $D_{1}(2430)$, has been found to appear at mass 2397 MeV even though the pion mass is larger than the corresponding physical mass, indicating that the pole mass could be lower. Updated calculations, with contributions from channels like $D\pi\pi$, are expected to be determined in future Lang and Wilson (2022), which can shed more light on this topic. As we will discuss, adding coupled channels like $D\rho$ can be important to better determine the properties of the $D_{1}$ states.
Several other attempts have been made to simultaneously describe the properties of the $D_{1}$ states Di Pierro and Eichten (2001); Bardeen et al. (2003); Colangelo et al. (2004); Mehen and Springer (2005); Ni et al. (2022); Kolomeitsev and Lutz (2004); Guo et al. (2007); Gamermann and Oset (2007); Malabarba et al. (2023); Coito et al. (2011); Burns (2014) but there does not seem to arise a consensus on the topic. In such a situation, where a clear manifestation of the states is not seen in the data and relevant information comes from a chi-square fit made to the data, while different models present different results, an alternative tool to test the dispersed findings is needed. We investigate if femtoscopic studies can play such a role.
@@
To carry out such a quest, we extend the framework developed in Refs. Gamermann and Oset (2007); Malabarba et al. (2023) by adding a quark-model pole to the lowest order $D^{*}\pi$ amplitude. We recall that a narrow pole is found to be generated from meson-meson coupled channel dynamics in the former works, which couples mainly to the $D\rho$ channel. The parameters used in Ref. Gamermann and Oset (2007) lead to the formation of such a pole around 2526 MeV. The parameters were fine-tuned in a later work Malabarba et al. (2023) and additional interaction diagrams (box diagrams, based on pion exchange) were added to the lowest order amplitude. As a consequence, a state was found whose mass and width were in excellent agreement with those of $D_{1}(2420)$. It’s worth emphasizing that the aforementioned pole couples strongly to $D\rho$ but very weakly to $D^{*}\pi$, which naturally explains its narrow width that is in contrast with the one of $D_{1}(2430)$. Next we should add that yet another pole, coupling mostly to $D^{*}\pi$, appears in the complex plane in the formalism of Ref. Malabarba et al. (2023), around $612222-i$ 61 MeV  [^5] The position of the corresponding pole in Ref. Gamermann and Oset (2007) was $115.682311.24-i$ 115.68 MeV, which is in better agreement with the properties of $D_{1}(2430)$. However, recall that the narrow pole found in Ref. Gamermann and Oset (2007) appears at $\sim$2526 MeV which is far from the mass of $D_{1}(2420)$.. This latter pole is not in good agreement with the properties of $D_{1}(2430)$. To improve this accordance, we add a bare quark-model pole to the lowest order $D^{*}\pi$ amplitude of Ref. Malabarba et al. (2023).
@@
On solving Bethe-Salpeter equations with such kernels, a broad state, as well as a narrow state, are found in the resulting amplitudes. The narrow state related to $D_{1}(2420)$ is as found in Ref. Malabarba et al. (2023). The broad state has properties closer to those of $D_{1}(2430)$. To further test the reliability of the model we calculate scattering lengths and compare the values for the $D^{*}\pi$ channel with those available from the lattice QCD calculations Mohler et al. (2013) done with a pion mass of 266 MeV. In the former work Mohler et al. (2013), the scattering length for $D\pi$ has also been obtained and a value very similar to the one for $D^{*}\pi$ is found. Such findings can indeed be understood in terms of heavy quark symmetry. In Refs. Liu et al. (2013); Guo et al. (2019), $D\pi$ and other systems are investigated in the lattice QCD formulation, the results of which are used to fix the parameters of the next-to-leading order term of the chiral Lagrangian and the $D\pi$ scattering length is obtained with the physical pion mass. Former values are in agreement with those obtained within unitarized chiral perturbation theory and heavy quark symmetry Guo et al. (2009); Abreu et al. (2011); Geng et al. (2010). Preliminary values for the $D\pi$ scattering length have also been determined by the ALICE Collaboration Alice (see also Refs. Grosa ; Battistini ; Fabbietti et al. ), though, enigmatically, the value determined seems to be different to the results of all the aforementioned works Liu et al. (2013); Guo et al. (2019, 2009); Abreu et al. (2011); Geng et al. (2010). Invoking arguments of heavy quark symmetry, one would expect that the scattering length would be similar for $D^{*}\pi$ and $D\pi$. The question which arises then is if the $D^{*}\pi$ scattering length should be expected to be similar to Refs. Liu et al. (2013); Guo et al. (2019, 2009); Abreu et al. (2011) or Ref. Alice . We contemplate both possibilities in the present manuscript. We find that small changes in the properties of the bare quark-model pole, added to the kernel, can produce a scattering length compatible with the one obtained for $D\pi$ in Refs. Liu et al. (2013); Guo et al. (2019, 2009); Abreu et al. (2011) or with Refs. Alice ; Grosa . Such changes affect the properties of the broader state only while the narrow state remains almost unaltered. Considering both parameterizations we determine the correlation functions for the different coupled systems involved and find that the results for the $D^{*}\pi$ channel are different in the two cases. We present results for different source sizes and discuss if such a tool can be used to decipher the properties of the two lowest-lying $D_{1}$ mesons.
@@
The manuscript is organized as follows. We first discuss the formalism used to determine the amplitudes for different meson-meson systems and present the convention followed to evaluate the scattering length. We give a summary of the scattering lengths obtained within different works, which leads to the consideration of two different parameterizations. We then show the amplitudes, discuss the properties of the resulting poles, and compare the scattering lengths obtained with other works. In a subsequent section, we present the details of the calculations of the correlation function. Finally, we summarize the conclusions of the work.
@@ II.1 Formalism to solve the scattering equations
Table 1: Values of the $C_{ij}$ constants appearing in Eq. (2) for different processes in the isospin 1/2 configuration. Here $\gamma=\left(\frac{m_{L}}{m_{H}}\right)^{2}$, where $m_{L}=800$ MeV and $m_{H}=2050$ MeV are average masses of the light and heavy vector-mesons, as defined in Refs. Gamermann and Oset (2007). In the case of isospin 3/2, we have two channels: $D^{*}\pi$ and $D\rho$ with the diagonal values of $C_{ij}$ being 1 and the non-diagonal being $\gamma$.
@@
with $P$ representing the total four-momentum of the system and $q$ the four-momentum of one of the mesons, $M_{k}$ ($m_{k}$) standing for the mass of the vector (pseudoscalar) meson in the $k$th channel, $a_{k}$ and $\mu$ denoting a subtraction constant needed to regularize the divergent nature of the $G$-function and the regularization scale respectively. The values for the regularization parameters are taken to be Malabarba et al. (2023) $\mu=1500$, $a=-1.45$, and we consider the convolution of $G$-function over the finite widths of $\rho$ and $K^{*}$. The resulting amplitudes show the presence of a state with mass ${}\sim 2428$ MeV and width $\sim 33$ MeV on the real axis, as in Ref. Malabarba et al. (2023). This latter state couples strongly to the $D\rho$ channel and weakly to the $D^{*}\pi$ channel, and its properties are in excellent agreement with those of $D_{1}(2420)$. There appears another pole at $612220-i$ 61 MeV, which couples mostly to $D^{*}\pi$ but which cannot be related to $D_{1}(2430)$ (see Fig. 1).
@@
Notice that the half-width of the narrow pole is around 5 MeV, in the complex plane, but the width on the real axis becomes 33 MeV on consideration of the finite widths of the $\rho$ and $K^{*}$ mesons. The width of the lower energy pole also increases, to  130 MeV, but still remains too small for the purpose of its association with $D_{1}(2430)$. Besides the mass also remains too low. We show the $D^{*}\pi$ and $D\rho$ squared amplitudes, on the real axis, as solid lines in Fig. 2.
Figure 2: Projection of the poles shown in Fig. 1 on the real axis, shown by depicting the squared amplitudes for the $D^{*}\pi$ and $D\rho$ channels. The solid lines are the results of solving the scattering equation with ten coupled channels (given in Table 1), while the dashed lines result from considering only the first four channels, which are $D^{*}\pi$, $D\rho$, $\bar{K}D^{*}_{s}$ and $D_{s}\bar{K}^{*}$. The two results look very similar, except for a $\eta D^{*}$ cusp effect seen around 2556 MeV in the $D^{*}\pi$ amplitude, shown as a solid line.
The results shown as dashed lines, Fig. 2, correspond to those obtained by considering only the first four channels of Table 1. It can be seen the results are almost unchanged, indicating that the most relevant channels to study the mentioned $D_{1}$ states are $D^{*}\pi$, $D\rho$, $\bar{K}D^{*}_{s}$ and $D_{s}\bar{K}^{*}$. The non-compatibility between the properties of $D_{1}(2430)$ and the lower energy pole seen in Fig. 1, whose effect on the real axis is shown in the left panel of Fig. 2, shows that something is missing in the model. We must recall at this point that the degrees of freedom considered in our model, so far, are hadrons and other inputs could be necessary to better describe the properties of $D_{1}(2420)$ and $D_{1}(2430)$ simultaneously.
For this purpose, as a next step, we try adding a bare quark-model pole to the lowest order amplitude for the $D^{*}\pi$ channel
@@
As we shall show next, the ansatz works well, and considering values of $M_{QM}$ from Refs. Godfrey and Isgur (1985); Ferretti and Santopinto (2018); Abreu et al. (2019) leads to the appearance of two poles in the amplitudes whose properties are in good accordance with the $D_{1}$ states of interest. The quark-model pole moves in the complex plane, as a consequence of non-perturbative meson-meson interactions, and becomes virtual, while the lower energy pole shown in Fig. 1 attains a larger mass and width. The reflection of such a pole shows properties in agreement with that of $D_{1}(2430)$.

### v2

# Can femtoscopic correlation function shed light on the nature of the lightest charm axial mesons?

@@ Abstract
We present a coupled channel treatment of meson-meson dynamics, for systems with spin-parity $1^{+}$, and determine the corresponding amplitudes by solving the Bethe-Salpeter equations, which lead to the generation of two axial resonances when mesons are considered as the degrees of freedom in the model. One of them is narrow and has properties in good agreement with those of $D_{1}(2420)$. The other pole is wider, but its real and imaginary parts do not match well with the mass and width of $D_{1}(2430)$. The situation improves when a bare quark-model state is included, indicating that the dynamics at the quark level as well as among hadrons can describe the two states simultaneously. Further, we discuss that there exists a divergence in the value of the $D^{*}\pi$ scattering length determined through data coming from lattice QCD calculations and from heavy ion collisions. Such different values can be accommodated in the model by making small changes in the parameters while producing two poles having properties compatible with the two lightest $D_{1}$ states. With these results, we proceed to calculate the correlation function for the $D^{*+(0)}\pi^{0(+)}$ system for different sizes of the source. We discuss which scenarios can be useful to shed some light on the issue.
@@ I Introduction
Femtoscopic correlation functions have emerged as a powerful alternative tool to understand the nature of hadrons in recent times (to see the articles published in the current year, for example, see Refs. Acharya et al. (2023a, b, c) or see the review of Ref. Fabbietti et al. (2021)), though the idea is not new Lisa et al. (2005). In the present manuscript, we analyze if the same idea can be implemented to better determine the nature of the two lightest $D_{1}$ mesons, $D_{1}(2420)$ and $D_{1}(2430)$, for which the average values of mass and width, as given by the Particle Data Group (PDG) Workman et al. (2022), are:
@@
The evidence for the existence of two $D_{1}$ states, with the aforementioned properties, comes from a fit made to the experimental data on the $D^{*}\pi$ invariant mass distribution Aaij et al. (2015); Abe et al. (2004), where the $S$- and $D$-wave amplitudes are associated with $D_{1}(2430)$ and $D_{1}(2420)$, respectively. From the theory side, two $c\bar{d}$, $c\bar{u}$ states were predicted, as expected from heavy quark symmetry, with the spin-parity $J^{\pi}=1^{+}$, in a quark model based on one-gluon-exchange-plus-linear-confinement potential Godfrey and Isgur (1985). Such states were found to arise from the $1^{1}P_{1}$ and $1^{3}P_{1}$ interactions, with masses 2440 and 2490 MeV, respectively, Godfrey and Isgur (1985). An admixture of the states is associated with the physical mesons, with a mixing angle of $-41^{0}$ Godfrey and Isgur (1985) (values of 35.30 or $-54.7^{0}$ for the mixing angle have been suggested based on the heavy quark symmetry Barnes et al. (2003)). Though the values of the masses determined in different quark models are not very different from those listed in Eq. (1), a difference in the width of the order of factor 10 cannot always be explained. Here, we must clarify that we are referring to models which try to explain the mass and width from the same dynamics. Indeed, arguments based on heavy quark symmetry have been presented by Manohar and Wise Manohar and Wise (2000) to explain that the ratio of the widths of the $D_{1}$ states could be of the order of factor 10, once the masses of the states are assumed and the quantum numbers $1^{1}P_{1}$ and $1^{3}P_{1}$ are attributed to $D_{1}(2430)$ and $D_{1}(2420)$, respectively (without considering a mixing between the two states). Indeed, as shown in Refs. Ferretti and Santopinto (2018); Cahn and Jackson (2003), a consideration of a mixing of the two quark model states through spin-orbit interactions leads to a ratio of widths of the two states to be far less than that found in the experimental data. In yet another work Zhou and Xiao (2011), it is shown that the ratio of the widths of the unmixed poles is around 2.5, which is not compatible with Eq. (1). The same work shows, however, that a mixing of the two states through the consideration of hadron loops can describe the widths of the two $D_{1}$ states and suggests that such a finding is compatible with the mixing angle proposed in Ref. Barnes et al. (2003). A similar conclusion has been reached in Ref. Ni et al. (2022). Further, a correction of about $-(5.7\pm 2.9)^{0}$ is proposed in Ref. Wang et al. (2018). In another study Godfrey (2005), predictions of radiative decay widths have been made to test the value of the mixing angle based on the heavy quark symmetry. Thus, it seems that different types of mixing of the $1^{1}P_{1}$ and $1^{3}P_{1}$ states lead to similar masses but different values of the widths of the $D_{1}$ states and that the consideration of hadron loops or meson clouds is useful in better describing the properties of the lowest-lying $D_{1}$ states.
On the other hand, the findings of models based purely on hadron dynamics show that although two low-lying $D_{1}$ states are also always found, their properties do not coincide with those found in experiments. Besides the difference in the properties of the states found in different models in comparison with those found experimentally, as we discuss below, there exists a deviation in the scattering length of $D^{*}\pi$ determined from data on lattice QCD Mohler et al. (2013) and from heavy ion collisions Acharya et al. (2024). Indeed information on observables related to the $D^{*}\pi$ channel is very relevant to understanding the nature of the $D_{1}$ states under discussion since it is the main decay channel for such states. Given the situation, one of the purposes of this manuscript is to propose a model that can describe the mass and width of the two lowest-lying $D_{1}$ states. Our attempt shows that an interplay of quark-hadron degrees of freedom can be useful in describing the aforementioned states (in line with the suggestions made in Ref. Zhou and Xiao (2011)). Yet another aim is to investigate if femtoscopic correlation functions related to channels and source sizes different from those considered in Ref. Acharya et al. (2024) can be useful in resolving the situation. We show that the extraction of data from smaller source sizes and for channels dominated by strong interactions (and not needing Coulomb interactions) can be used to settle the value of the $D^{*}\pi$ scattering length, which can contribute to a better understanding of the properties of the two lowest-lying $D_{1}$ states.
I.1 A short summary of model findings.
@@
Let us summarize the results of some of the works attempting to describe the dynamical origin of the mass and width of the lightest axial states with charm. For example, it was pointed out in Ref. Du et al. (2018) that all the low-lying positive-parity heavy open-flavor mesons can be understood as hadronic molecules. Such a claim is motivated by facts like the masses of the lightest open-charm mesons with strangeness are lower than their nonstrange counterparts. In Ref. Du et al. (2018), scattering equations were solved using kernels based on the unitarized chiral perturbation theory for heavy mesons and by fixing the free parameters to fit the scattering lengths determined from lattice QCD calculations. Within such a framework, two poles were found for $D_{1}$ mesons, with the mass and the half-width being $2247^{+5}_{-6},~{}107^{+11}_{-10}$ and $2555^{+47}_{-30},~{}203^{+8}_{-9}$, where the lower one was associated with $D_{1}(2430)$. It is argued in the former work that the mass value, obtained from a Breit-Wigner fit, given by PDG for $D_{1}(2430)$ needs to be revised. A narrow pole, to be associated with $D_{1}(2420)$, however, is not found. Curiously, a very different quark-model calculation Abreu et al. (2019) obtained mass values for the $D_{1}$ states which are very similar to those of Ref. Du et al. (2018). Information on $D_{1}$ mesons is also available from ${}^{2S+1}l_{J}={}^{3}S_{1},~{}{}^{3}D_{1}$ amplitudes determined for the $D^{*}\pi$ system with lattice QCD calculations Lang and Wilson (2022). A broad and a narrow state is found, respectively, in the former and the latter amplitudes. It is worth mentioning that the broad ${}^{3}S_{1}$ state, which is related to $D_{1}(2430)$, has been found to appear at mass 2397 MeV even though the pion mass is larger than the corresponding physical mass, indicating that the physical pole mass could be lower. Updated calculations, with contributions from channels like $D\pi\pi$, are expected to be determined in future Lang and Wilson (2022). As we will discuss, in fact adding coupled channels like $D\rho$ can be important to better understand the properties of the $D_{1}$ states.
It is worth mentioning that several other attempts have been made to simultaneously describe the properties of the $D_{1}$ states Di Pierro and Eichten (2001); Colangelo et al. (2004); Kolomeitsev and Lutz (2004); Guo et al. (2007); Gamermann and Oset (2007); Malabarba et al. (2023); Coito et al. (2011); Burns (2014), claiming different attributions to their dynamical origin
I.2 Our idea
@@
The main purpose of our work is to determine the femtoscopic correlation functions for the $D^{*+(0)}\pi^{0(+)}$ systems, which are free from Coulomb interactions. These systems are different from those investigated in Ref. Acharya et al. (2024). The idea is also to calculate the correlation function for different sizes of the source and study if there exist more ideal conditions to determine the nature of the $D^{*}\pi$ strong interaction. To do this we need the amplitudes of the $D^{*}\pi$ system which carry the information on the $D_{1}$ states. We proceed by extending the framework developed in Refs. Gamermann and Oset (2007); Malabarba et al. (2023) by adding a quark-model pole to the lowest order $D^{*}\pi$ amplitude. As shown in Ref. Malabarba et al. (2023) the dynamics of $D^{*}\pi$ and coupled channels leads to the formation of a state whose mass and width are in excellent agreement with those of $D_{1}(2420)$. It is worth emphasizing that the aforementioned pole couples strongly to $D\rho$ but very weakly to $D^{*}\pi$, which naturally explains its narrow width that is in contrast with the one of $D_{1}(2430)$. Next, we should add that yet another pole, coupling mostly to $D^{*}\pi$, appears in the complex plane in the formalism of Ref. Malabarba et al. (2023), around $612222-i$ 61 MeV  [^5] The position of the corresponding pole in Ref. Gamermann and Oset (2007) was $115.682311.24-i$ 115.68 MeV, which is in better agreement with the properties of $D_{1}(2430)$. However, recall that the narrow pole found in Ref. Gamermann and Oset (2007) appears at $\sim$2526 MeV which is far from the mass of $D_{1}(2420)$.. This latter pole is not in good agreement with the properties of $D_{1}(2430)$. To improve this accordance, we add a bare quark-model pole to the lowest order $D^{*}\pi$ amplitude of Ref. Malabarba et al. (2023).
@@
On solving Bethe-Salpeter equations with such kernels, a broad state, as well as a narrow state, are found in the resulting amplitudes. The narrow state related to $D_{1}(2420)$ is as found in Ref. Malabarba et al. (2023). The broad state has properties closer to those of $D_{1}(2430)$. To further compare the results of the model with the known data we calculate the $D^{*}\pi$ scattering length. We recall that scattering length values for the $D\pi$ and $D^{*}\pi$ channels are determined in Ref. Mohler et al. (2013) within the lattice QCD framework, using a pion mass of 266 MeV. Also, in Refs. Liu et al. (2013); Guo et al. (2019), $D\pi$ and other systems are investigated in the lattice QCD formulation, the results of which are used to fix the parameters of the next-to-leading order term of the chiral Lagrangian and the $D\pi$ scattering length is obtained with the physical pion mass. The resulting values are found to agree with those obtained within unitarized chiral perturbation theory and heavy quark symmetry Guo et al. (2009); Abreu et al. (2011); Geng et al. (2010). Values of the $D\pi$ and $D^{*}\pi$ scattering lengths have also been determined by the ALICE collaboration Acharya et al. (2024) (see also Refs. Alice ; Grosa ; Battistini ; Fabbietti et al. ), though, enigmatically, the value determined seems to be different to the results of all the aforementioned works Liu et al. (2013); Guo et al. (2019, 2009); Abreu et al. (2011); Geng et al. (2010). The question that arises is if the $D\pi$ and $D^{*}\pi$ scattering lengths should be as obtained in Refs. Liu et al. (2013); Guo et al. (2019, 2009); Abreu et al. (2011) or as in Ref. Acharya et al. (2024). It is the purpose of our paper to investigate if femtoscopic studies of systems or source sizes different from those considered in Ref. Acharya et al. (2024) can be used to resolve this matter. Hence we contemplate two possibilities in our model which can produce different scattering lengths and calculate the correlation function for $D^{*0(+)}\pi^{+(0)}$, for different values of the source size. We present results for different source sizes and discuss if such a tool can be used to resolve the issue and discuss the implications of the two scenarios on the properties of the two lowest-lying $D_{1}$ mesons.
@@
The article is organized as follows. We first discuss the formalism used to determine the amplitudes for different meson-meson systems and present the convention followed to evaluate the scattering length. We give a summary of the scattering lengths obtained within different works, which leads to the consideration of two different parametrizations. We then show the amplitudes, discuss the properties of the resulting poles, and compare the scattering lengths obtained with other works. In a subsequent section, we present the details of the calculations of the correlation function. Finally, we summarize the conclusions of the work.
@@ II.1 Formalism to solve the scattering equations
Table 1: Values of the $C_{ij}$ constants appearing in Eq. (2) for different processes in the isospin 1/2 configuration. Here $\gamma=\left(\frac{m_{L}}{m_{H}}\right)^{2}$, where $m_{L}=800$ MeV and $m_{H}=2050$ MeV are average masses of the light and heavy vector mesons, as defined in Refs. Gamermann and Oset (2007). In the case of isospin 3/2, we have two channels: $D^{*}\pi$ and $D\rho$ with the diagonal values of $C_{ij}$ being 1 and the nondiagonal being $\gamma$.
@@
with $P$ representing the total four-momentum of the system and $q$ the four-momentum of one of the mesons, $M_{k}$ ($m_{k}$) standing for the mass of the vector (pseudoscalar) meson in the $k$th channel, $a_{k}$ and $\mu$ denoting a subtraction constant needed to regularize the divergent nature of the $G$-function and the regularization scale, respectively. The values for the regularization parameters are taken to be Malabarba et al. (2023) $\mu=1500$, $a=-1.45$, and we consider the convolution of $G$-function over the finite widths of $\rho$ and $K^{*}$. The resulting amplitudes show the presence of a state with mass ${}\sim 2428$ MeV and width $\sim 33$ MeV on the real axis, as in Ref. Malabarba et al. (2023). This latter state couples strongly to the $D\rho$ channel and weakly to the $D^{*}\pi$ channel, and its properties are in excellent agreement with those of $D_{1}(2420)$. There appears another pole at $612220-i$ 61 MeV, which couples mostly to $D^{*}\pi$ but which cannot be related to $D_{1}(2430)$ (see Fig. 1).
@@
Notice that the half-width of the narrow pole is around 5 MeV, in the complex plane, but the width on the real axis becomes 33 MeV on consideration of the finite widths of the $\rho$ and $K^{*}$ mesons. The width of the lower energy pole also increases to  130 MeV, but it still remains too small for the purpose of its association with $D_{1}(2430)$. Besides the mass also remains too low. We show the $D^{*}\pi$ and $D\rho$ squared amplitudes, on the real axis, as solid lines in Fig. 2.
Figure 2: Projection of the poles shown in Fig. 1 on the real axis, shown by depicting the squared amplitudes for the $D^{*}\pi$ and $D\rho$ channels. The solid lines are the results of solving the scattering equation with ten coupled channels (given in Table 1), while the dashed lines result from considering only the first four channels, which are $D^{*}\pi$, $D\rho$, $\bar{K}D^{*}_{s}$, and $D_{s}\bar{K}^{*}$. The two results look very similar, except for a $\eta D^{*}$ cusp effect seen around 2556 MeV in the $D^{*}\pi$ amplitude, shown as a solid line.
The results shown as dashed lines, Fig. 2, correspond to those obtained by considering only the first four channels of Table 1. It can be seen the results are almost unchanged, indicating that the most relevant channels to study the mentioned $D_{1}$ states are $D^{*}\pi$, $D\rho$, $\bar{K}D^{*}_{s}$, and $D_{s}\bar{K}^{*}$. The noncompatibility between the properties of $D_{1}(2430)$ and the lower energy pole seen in Fig. 1, whose effect on the real axis is shown in the left panel of Fig. 2, shows that something is missing in the model. We must recall at this point that the degrees of freedom considered in our model, so far, are hadrons and other inputs could be necessary to better describe the properties of $D_{1}(2420)$ and $D_{1}(2430)$ simultaneously.
For this purpose, as a next step, we try adding a bare quark-model pole to the lowest order amplitude for the $D^{*}\pi$ channel, following Refs. Nieves and Pavao (2020); Cincioglu et al. (2016); Albaladejo et al. (2017, 2018); Geng et al. (2007); Martínez Torres et al. (2015), as [^6] The potential in Eq. (6) should be considered as an effective interaction added to one of the channels mimicking the effect of adding a term to each channel. Such an ansatz is followed to minimize the number of free parameters
@@
Before proceeding further we must clarify why we add one quark-model pole when two of them appear in Refs. Godfrey and Isgur (1985); Ferretti and Santopinto (2018); Abreu et al. (2019). The reason is that one of the states generated by meson-meson dynamics, in our formalism (which is based on Refs. Gamermann and Oset (2007); Malabarba et al. (2023)), is in very good agreement with the properties of $D_{1}(2420)$. The implication of such a finding is that the contribution of the quark-model component to the $D_{1}(2420)$ wave function must be small, such as to keep the properties of the pole related to $D_{1}(2420)$ mostly unchanged. It should also be mentioned that the regularization parameter used to solve the Bethe-Salpeter equation can embed contributions from other types of dynamics not explicitly considered in the model (see discussions in Refs. Hyodo et al. (2012); Martinez Torres et al. (2012); Aceti and Oset (2012)).